### v2

@@ Abstract
This paper presents a comparative analysis evaluating the accuracy of Large Language Models (LLMs) against traditional macro time series forecasting approaches. In recent times, LLMs have surged in popularity for forecasting due to their ability to capture intricate patterns in data and quickly adapt across very different domains. However, their effectiveness in forecasting macroeconomic time series data compared to conventional methods remains an area of interest. To address this, we conduct a rigorous evaluation of LLMs against traditional macro forecasting methods, using as common ground the FRED-MD database. Our findings provide valuable insights into the strengths and limitations of LLMs in forecasting macroeconomic time series, shedding light on their applicability in real-world scenarios.
@@ 1 Introduction
Our first contribution in this paper is a thorough investigation on how TSLMs perform in predicting macroeconomic time series. A second contribution is a detailed comparison of their performance versus state-of-the-art time series methods such as Bayesian Vector Autoregressions (BVARs) and Factor Models. In these endeavours, we focus on forecasting the variables contained in the FRED-MD dataset, a comprehensive repository of more than one hundred monthly macroeconomic variables curated by the Federal Reserve Economic Data (FRED) system. [^3] See https://research.stlouisfed.org/econ/mccracken/fred-databases/. This dataset encompasses a diverse array of economic indicators, and it is considered a benchmark in macroeconomic forecasting. A non-exhaustive lists of contributions using this data set to study the performance of alternative models in macroeconomic forecasting includes Stock and Watson (2006), Banbura et al. (2010), Carriero et al. (2016), Carriero et al. (2019), Korobilis and Pettenuzzo (2019), Koop et al. (2019), Chan (2022).
@@
As noted in Faria-e Castro and Leibovici (2024), there are some characteristics of LLM models that makes it problematic to use them in the pseudo out-of-sample forecasting exercise that is typically performed in the empirical macroeconomics literature. One major challenge is that LLMs are pre-trained by developers on datasets of their choice, giving the researcher little to no control over the training data. For example, of the LLMs considered in this work, three out of five of them (including the best performing one, Moirai) list in their training data a large subset of the series we set out to forecast in this paper. To further complicate matters, since the training data lacks timestamps it’s hard to retrain the model up to a certain date, making it difficult to reproduce the type of real-time analysis that an economic forecaster is mostly concerned with.
Solving these issues by retraining LLMs on subsets of data and adding data to the training set is usually not possible due to high hardware requirements. There are some exceptions, such as for example LagLlama, which could now be fine-tuned in real time at relatively lower costs. However, this paper will focus on “zero-shot” forecasting and such fine-tuning is beyond its scope. [^4] By “zero-shot”, we mean that the model is first trained on a large amount of data, and is then directly applied to solve a new task without re-training, or fine-tuning, its parameters on the new dataset. Moreover, even including the dataset of interest in real time would not remove the issue of a contaminated training set, i.e. including information that the forecaster would not have access to in real-time.
@@
Our empirical application is focused on forecasting the FRED monthly database using data ranging from 1960 to 2023. When working with traditional time series methods, we consider models of different cross-sectional dimensions, in order to understand the role that the information set has on the forecasting performance. The picture emerging from our analysis is one in which only two of the five TSLMs we consider are competitive against a simple AR benchmark (Salesforce’s Moirai and Google’s TimesFM). Moreover, when these more competitive TSLMs are stacked against Bayesian VARs and Factor Models, their forecasting performance is broadly comparable, if not slightly inferior. In particular, we find that the forecasting gains achieved by the econometric models tend to be more stable while TSLMs can perform very well for a handful of series but also show less reliability at times, as they seem to be more prone to generating the occasional unreasonable forecast. [^5] This result may not appear to some as totally surprising. When working with textual LLMs, we are sometimes exposed to “Model hallucinations”, i.e. the tendency of the model to generate content that is irrelevant, made-up, or inconsistent with the input data. On the other hand, TSLMs seem to work relatively better in the post-COVID era (with the important caveat that the training set of these models does include information from the pandemic and post-pandemic period).
@@
There are a few other papers that are looking at LLM in the macroeconomics and finance literature, and it is reasonable to expect that many more will follow. For example Bybee (2023) feed into an LLM model, OpenAI’s GPT-3.5, a historical sample of news articles from The Wall Street Journal (WSJ) and ask the model to predict various financial and macroeconomic quantities, which they then aggregate into a time-series of monthly and quarterly expectations that are compared to a variety of existing survey forecasts. Chen et al. (2022) feed global news text data from Thomson Reuters Real-time News Feed (RTRS) and Third Party Archive (3PTY) in 13 different languages into pre-trained LLMs, namely BERT (developed by Google), RoBERTa (by Meta), and OPT (by Meta), and use the resulting sentiment scores to predict firm-level daily returns. Kim et al. (2024) feed standardized and anonymous financial statements to GPT4 and instruct the model to analyze them to determine the direction of future earnings. They find that LLMs outperforms financial analysts in their ability to predict earnings changes. These approaches focus on leveraging the typical LLMs abilities in natural language processing, while this paper is focused on the recent efforts of training LLMs specifically for time-series forecasting. The closest work we are aware of in the macroeconomics literature is Faria-e Castro and Leibovici (2024), which looks at time series LLMs, but only focuses on one LLM and one target variable (inflation). In contrast to this, in this paper we take on the ambition goal of looking at the performance of a variety of LLMs, comparing their accuracy against state of the art macroeconomic forecasting approaches, across a large set of macroeconomic indicators.
@@
The paper is organized as follows. Section 2 introduces LLMs and TSLMs and provides a discussion on how they are being leveraged to perform time series forecasting. Section 3 describes the econometric models we consider as state of the art benchmarks in our exercise. Section 4 describes the dataset and the design of the out-of-sample pseudo real time forecasting exercise and Section 5 presents our empirical results. Finally, Section 6 provides some concluding remarks and directions for further research. An Appendix provides technical details on the models considered and the data transformations used in the empirical analysis.
@@ 2.1 Time Series Language Models (TSLMs)
Inspired by these successes, researchers are now exploring the usefulness of LLMs in other settings. Along these lines, Lu et al. (2022) find that LLMs trained exclusively on textual data can effectively solve sequence modeling tasks in other modalities, such as numerical computation, vision, and protein fold prediction. Among these extensions, we are witnessing the emergence of Time Series Language Models (TSLM), bridging the gap between LLMs original text data training and the numerical nature of time series data. These models are sometimes referred to in the literature as Time Series Foundational Models (TSFM). TSLMs are now being used to accomplish a variety of time series related tasks, ranging from prediction and classification to anomaly detection and data imputation. Notable examples of this new wave of model development include Rasul et al. (2024), Goswami et al. (2024), Ekambaram et al. (2024), Garza and Mergenthaler-Canseco (2023), Das et al. (2024), and Ansari et al. (2024). At the core of this very recent advancements is a simple but fundamental point. While there are in principle significant differences between a textual LLM, whose main task is essentially to predict the next word, and a time series forecasting model, whose main objective is to predict the next value of one or multiple time series jointly, inherently these two tasks are deeply intertwined. They both aim at learning the sequential structure of the data from the past and present history of some features and with that attempt to predict a sequence of future outcomes. In the case of a TSLM, the pre-training is done on a possibly very large set of time series, $\bm{X}_{1:T}=\left(x_{1,1:T},...,x_{N,1:T}\right)$, and yields a mapping function $f_{\theta}$. Then, given the current and past values of of a time series of interest $y$ (not necessarily in the training set), the TSLM attempts to predict its future $h$ values $y_{T+1:T+h}$ using [^6] To keep the notation simple, we have assumed here a balanced set of features, i.e. all available over the same time period. If that is not the case, then we could instead write this as $\bm{X}_{1:T}=\left(x_{1,\underline{t}_{1}:T_{1}},...,x_{N,\underline{t}_{N}:T_{N}}\right)$, where $\underline{t}_{i}$ and $T_{i}$ would denote, respectively, the first and last observation of feature $i$, $i\in\left\{1,...,N\right\}$.
$f_{\theta}\left(\left.y_{T+1:T+h}\right|y_{1:T};\bm{X}_{1:T},\theta\right).$ (1)
@@ 2.2 Components of Time Series Language Models
The most recent versions of TSLMs have evolved from this initial approach and now build on quantization-based methods, i.e. a process through which the numerical data is converted into discrete representations first, before becoming an input to the training of the LLM. We now turn to describing the basic building blocks of a TSLM, drawing whenever possible parallels to their textual counterparts for clarity. The specific choices and configurations of these building blocks define the unique characteristics of the various models.
@@ 2.2.3 Time Series Augmentation
2.2.4 Pre-training and Fine-tuning
Pre-training refers to the process of feeding to the TSLMs a large amount of time series data so that the model can learn general temporal patterns and relationships and with that perform well across a wide spectrum of tasks. The time series data fed into the TSLM can be from various domains and span different frequencies and, as previously discussed, each time series will first be scaled and tokenized before any training occurs. Once the pre-training phase is completed, the TSLMs will have built a foundational knowledge base from a very general corpus of time series data, and with that can be used to perform zero-shot forecasting in a new domain.
Pre-training relies on self-supervised techniques that create tasks for the model so that it can learn directly from the input data without requiring human labeled examples. The success rate in these tasks can be measured in various ways, such as minimizing MSE or accurately predicting the quantized bins or the quantiles of the training data. As a concrete example, LagLlama pre-training looks for the model parameters that maximize the likelihood of the observed future values, under the assumption that these follow a Student-t distribution.
Fine-tuning is a training process that is used to adapt a pre-trained TSLM, containing general foundational knowledge, to a specific task such as forecasting, or classification. Fine-tuning updates existing knowledge (parameters) of TSLM to better handle unique characteristics of the task-specific data. This can potentially improve the performance of a TSLM compared to zero-shot predictions for the given task. Fine-tuning can also be used to continually learn the model parameters and incorporate fresh data, improving TSLM’s forecasting accuracy over time. For example, in macroeconomic forecasting, we can use fine-tuning to incorporate the data from the current quarter to update the parameters of the TSLM.
@@ 2.3 Relationship with Deep Learning
Before moving on to describe the existing TSLMs, it is worth briefly discussing the difference between Deep Learning (DL) models (see for example Torres et al. (2021), Lim and Zohren (2021) and Wen et al. (2022)) and TSLMs. DL models, share many similarities with traditional time-series methods, requiring problem specific training data to learn model parameters and make forecasts. However, they differ the traditional methods in a number of ways – using non-linear architectures, large numbers of learnable parameters, and requiring much larger training data than traditional time series methods. TSLMs, on the other hand, are pre-trained on large number of time series datasets. As with FMs, this pre-training allows them to perform “zero-shot” forecasting. This plug-and-play approach makes TSLMs particularly useful and powerful. Furthermore, pre-training allows TSLMs to require much less domain-specific knowledge data compared to deep learning models, and if needed can be fine-tuned on specific datasets.
@@ 2.4 State-of-the-art TSFMs
We now briefly describe the TSFM models considered in this paper. [^12] In addition to the models listed in this section, we have also experimented with MOMENT (Goswami et al., 2024), a family of time series transformer models that is pre-trained on masked time series prediction task using patching. However this model is still in its very early stages and we have found its performance to be lagging significantly below the other TSFMs we have considered. We therefore did not include it in the comparison below.
@@
(Rasul et al., 2024) LagLlama is a probabilistic foundation model that extends the Llama 2 architecture (Touvron et al., 2023) to time series data. LagLlama follows a z-score normalization at a window level, and tokenizes the input time series by extracting lagged features from the past values of the time series at different time lags. This allows the model to learn how past observations influence future ones. LagLlama is pre-trained on 27 publicly available time series datasets from various domains. The model has about $\sim 2.5$M parameters.
@@
Table 1 lists a few additional details on the various TSLMs described in this section. In particular, it provides information on when the model was released, i.e. when the pre-training sample ends, the list of all the data domains used in the pre-training phase of the model, and the size of training sample. It also flags whether each TSLM is univariate or multivariate.
@@
Table 1: Training datasets for the various TSLMs considered in this paper. For each model, we provide a summary of when the model was pre-trained, the list of all the datasets used in the pre-training phase of the model, the size of training sample, and whether it is multivariate. Additional details and references on the various datasets can be found in the individual TSLMs’ technical papers. indicates that the training data include the Monash forecast repository (available at https://forecastingdata.org), and therefore includes a large part of the FRED-MD dataset.

### v3

@@ Abstract
This paper presents a comparative analysis evaluating the accuracy of Large Language Models (LLMs) against traditional macro time series forecasting approaches. In recent times, LLMs have gained in popularity for forecasting due to their ability to capture intricate patterns in data and quickly adapt across very different domains. However, their effectiveness in forecasting macroeconomic time series data compared to conventional methods remains an area of interest. To address this, we conducted a rigorous evaluation of LLMs against traditional macro-prediction methods, using as common ground the FRED-MD database. Our findings provide valuable insights into the strengths and limitations of LLMs in the prediction of macroeconomic time series, shedding light on their applicability in real world scenarios.
@@ 1 Introduction
Our first contribution in this paper is a thorough investigation on how TSLMs perform in predicting macroeconomic time series. A second contribution is a detailed comparison of their performance versus state-of-the-art time series methods such as Bayesian Vector Autoregressions (BVARs) and Factor Models as well as Bayesian Additive Regression Trees (BART), which has become a very popular nonparametric model for macroeconomic forecasting. In these endeavors, we focus on forecasting the variables contained in the FRED-MD dataset, a comprehensive repository of more than one hundred monthly macroeconomic variables curated by the Federal Reserve Economic Data (FRED) system. [^3] See https://research.stlouisfed.org/econ/mccracken/fred-databases/. This dataset encompasses a diverse array of economic indicators, and it is considered a benchmark in macroeconomic forecasting. A non-exhaustive lists of contributions using this data set to study the performance of alternative models in macroeconomic forecasting includes Stock and Watson (2006), Banbura et al. (2010), Carriero et al. (2016), Carriero et al. (2019), Korobilis and Pettenuzzo (2019), Koop et al. (2019), Chan (2022).
@@
As noted in Faria-e Castro and Leibovici (2024), there are some characteristics of LLM models that makes it problematic to use them in the pseudo out-of-sample forecasting exercise that is typically performed in the empirical macroeconomics literature. One major challenge is that LLMs are pretrained by developers on datasets of their choice, giving the researcher little to no control over the training data. For example, of the LLMs considered in this work, three out of five of them (including the best performing one, Moirai) list in their training data a large subset of the series we set out to forecast in this paper. To further complicate matters, since the training data lacks timestamps it’s hard to retrain the model up to a certain date, making it difficult to reproduce the type of real-time analysis that an economic forecaster is mostly concerned with. Solving these issues by retraining LLMs on subsets of data and adding data to the training set is usually not possible due to high hardware requirements. But having said that, it is important to recognize that even including the dataset of interest in real time would not remove the issue of a contaminated training set, i.e. including information that the forecaster would not have access to in real-time.
@@
Our empirical application is focused on forecasting the FRED monthly database using data ranging from 1960 to 2023. When working with traditional time series methods, we consider models of different cross-sectional dimensions, in order to understand the role that the information set has on the forecasting performance. The picture emerging from our analysis is one in which only two of the five TSLMs we consider are competitive against a simple AR benchmark (Salesforce’s Moirai and Google’s TimesFM). Moreover, when these more competitive TSLMs are stacked against Bayesian VARs and Factor Models, their forecasting performance is broadly comparable, if not slightly inferior. In particular, we find that the forecasting gains achieved by the econometric models tend to be more stable while TSLMs can perform very well for a handful of series but also show less reliability at times, as they seem to be more prone to generating the occasional unreasonable forecast. [^4] This result may not appear to some as totally surprising. When working with textual LLMs, we are sometimes exposed to “Model hallucinations”, i.e. the tendency of the model to generate content that is irrelevant, made-up, or inconsistent with the input data. On the other hand, TSLMs seem to work relatively better in the post-Covid-19 era (with the important caveat that the training set of these models does include information from the pandemic and post-pandemic period). While these results are based on using the TSLM models in a zero-shot forecasting manner, we also find that fine-tuning the TSLM models does not yield to significant improvements in their forecast accuracy. [^5] By “zero-shot”, we mean that the model is first trained on a large amount of data, and is then directly applied to solve a new task without re-training, or fine-tuning, its parameters on the new dataset.
@@
There are a few other papers that are looking at LLM in the macroeconomics and finance literature, and it is reasonable to expect that many more will follow. For example Bybee (2023) feed into an LLM model, OpenAI’s GPT-3.5, a historical sample of news articles from The Wall Street Journal (WSJ) and ask the model to predict various financial and macroeconomic quantities, which they then aggregate into a time-series of monthly and quarterly expectations that are compared to a variety of existing survey forecasts. Chen et al. (2022) feed global news text data from Thomson Reuters Real-time News Feed (RTRS) and Third Party Archive (3PTY) in 13 different languages into pretrained LLMs, namely BERT (developed by Google), RoBERTa (by Meta), and OPT (by Meta), and use the resulting sentiment scores to predict firm-level daily returns. Kim et al. (2024) feed standardized and anonymous financial statements to GPT4 and instruct the model to analyze them to determine the direction of future earnings. They find that LLMs outperforms financial analysts in their ability to predict earnings changes. These approaches focus on leveraging the typical LLMs abilities in natural language processing, while this paper is focused on the recent efforts of training LLMs specifically for time-series forecasting. Chen et al. (2025) study whether both DeepSeek and ChatGPT can be used to predict the stock market and find that a good news ratio extracted from prompting ChatGPT is positively correlated with contemporaneous market returns and significantly predicts subsequent returns up to six months out. The closest work we are aware of in the macroeconomics literature is Faria-e Castro and Leibovici (2024), which looks at time series LLMs, but only focuses on one LLM and one target variable (inflation). In contrast to this, in this paper we take on the ambition goal of looking at the performance of a variety of LLMs, comparing their accuracy against state of the art macroeconomic forecasting approaches, across a large set of macroeconomic indicators.
@@
The paper is organized as follows. Section 2 introduces LLMs and TSLMs and provides a discussion on how they are being leveraged to perform time series forecasting. Section 3 describes the econometric models we consider as state of the art benchmarks in our exercise. Section 4 describes the dataset and the design of the out-of-sample pseudo real time forecasting exercise and Section 5 presents our empirical results. Finally, Section 6 describes our fine-tuning process and the associated results, while Section 7 provides some concluding remarks and directions for further research. An Appendix provides technical details on the models considered and the data transformations used in the empirical analysis.
@@ 2.1 Time Series Language Models (TSLMs)
Inspired by these successes, researchers are now exploring the usefulness of LLMs in other settings. Along these lines, Lu et al. (2022) find that LLMs trained exclusively on textual data can effectively solve sequence modeling tasks in other modalities, such as numerical computation, vision, and protein fold prediction. Among these extensions, we are witnessing the emergence of Time Series Language Models (TSLM), bridging the gap between LLMs original text data training and the numerical nature of time series data. These models are sometimes referred to in the literature as Time Series Foundational Models (TSFM). TSLMs are now being used to accomplish a variety of time series related tasks, ranging from prediction and classification to anomaly detection and data imputation. Notable examples of this new wave of model development include Rasul et al. (2024), Goswami et al. (2024), Ekambaram et al. (2024), Garza and Mergenthaler-Canseco (2023), Das et al. (2024), and Ansari et al. (2024). At the core of this very recent advancements is a simple but fundamental point. While there are in principle significant differences between a textual LLM, whose main task is essentially to predict the next word, and a time series forecasting model, whose main objective is to predict the next value of one or multiple time series jointly, inherently these two tasks are deeply intertwined. They both aim at learning the sequential structure of the data from the past and present history of some features and with that attempt to predict a sequence of future outcomes. In the case of a TSLM, the pretraining is done on a possibly very large set of time series, $\bm{X}_{1:T}=\left(x_{1,1:T},...,x_{N,1:T}\right)$, and yields very large set of coefficients $\theta^{*}$ which describe a forecasting function $f_{\theta^{*}}$. Then, given $f_{\theta^{*}}$ and the current and past values of a time series of interest $y$ (not necessarily in the training set), the TSLM attempts to predict its future $h$ values $y_{T+1:T+h}$ using [^6] To keep the notation simple, we have assumed here a balanced set of features, i.e. all available over the same time period. If that is not the case, then we could instead write this as $\bm{X}_{1:T}=\left(x_{1,\underline{t}_{1}:T_{1}},...,x_{N,\underline{t}_{N}:T_{N}}\right)$, where $\underline{t}_{i}$ and $T_{i}$ would denote, respectively, the first and last observation of feature $i$, $i\in\left\{1,...,N\right\}$.
$\widehat{y}_{T+1:T+H|T}=f_{\theta^{*}}(\left.y_{1:T}\right)=f(\left.y_{1:T}|\bm{X}_{1:T},\theta^{*}\right).$ (1)
@@ 2.2 Components of Time Series Language Models
The most recent versions of TSLMs have evolved from this initial approach and now build on quantization-based methods, i.e. a process through which the numerical data is converted into discrete representations first, before becoming an input to the training of the LLM. We now turn to describing the basic building blocks of a TSLM, drawing whenever possible parallels to their textual counterparts for clarity. The specific choices and configurations of these building blocks define the unique characteristics of the various models. An overview of TSLM architecture is shown in Figure 1.
Figure 1: Components of a TSLM: The key components of a TSLM include tokenization, pretraining, and zero-shot forecasting. First, raw time series is tokenized into discrete tokens using techniques like scaling, quantization, and patching. Tokens are then used to pretrain a Transformer-based model that learns temporal dependencies and patterns. Finally, through zero-shot forecasting, the pretrained model can forecast in new and unseen time series domains. Note that different techniques at each stage–such as tokenization methods and model architectures–leads to different TSLM architectures.
@@ 2.2.3 Time Series Augmentation
2.2.4 Pretraining and Fine-tuning
Pretraining
Pretraining refers to the process of feeding to the TSLMs a large amount of time series data so that the model can learn general temporal patterns and relationships and with that perform well across a wide spectrum of tasks. The time series data fed into the TSLM can be from various domains and span different frequencies and, as previously discussed, each time series will first be scaled and tokenized before any training occurs. Once the pretraining phase is completed, the TSLMs will have built a foundational knowledge base from a very general corpus of time series data, and with that can be used to perform zero-shot forecasting in a new domain.
Pretraining relies on self-supervised techniques that create tasks for the model so that it can learn directly from the input data without requiring human labeled examples. The success rate in these tasks can be measured in various ways, such as minimizing MSE or accurately predicting the quantized bins or the quantiles of the training data. As a concrete example, LagLlama pretraining looks for the model parameters that maximize the likelihood of the observed future values, under the assumption that these follow a Student-t distribution.
Fine-tuning
Fine-tuning is a training process that is used to adapt a pretrained TSLM, containing general foundational knowledge, to a specific task such as forecasting, or classification. Fine-tuning updates existing knowledge (parameters) of TSLM to better handle unique characteristics of the task-specific data. This can potentially improve the performance of a TSLM compared to zero-shot predictions for the given task. Fine-tuning can also be used to continually learn the model parameters and incorporate fresh data, improving TSLM’s forecasting accuracy over time. For example, in macroeconomic forecasting, we can use fine-tuning to incorporate the data from the current quarter to update the parameters of the TSLM.
@@ 2.3 Relationship with Deep Learning
Before moving on to describe the existing TSLMs, it is worth briefly discussing the difference between Deep Learning (DL) models (see for example Torres et al. (2021), Lim and Zohren (2021) and Wen et al. (2022)) and TSLMs. DL models, share many similarities with traditional time-series methods, requiring problem specific training data to learn model parameters and make forecasts. However, they differ the traditional methods in a number of ways – using non-linear architectures, large numbers of learnable parameters, and requiring much larger training data than traditional time series methods. TSLMs, on the other hand, are pretrained on large number of time series datasets. As with FMs, this pretraining allows them to perform “zero-shot” forecasting. This plug-and-play approach makes TSLMs particularly useful and powerful. Furthermore, pretraining allows TSLMs to require much less domain-specific knowledge data compared to deep learning models, and if needed can be fine-tuned on specific datasets.
@@ 2.4 State-of-the-art TSFMs
We now briefly describe the TSFM models considered in this paper. [^12] In addition to the models listed in this section, we have also experimented with MOMENT (Goswami et al., 2024), a family of time series transformer models that is pretrained on masked time series prediction task using patching. However this model is still in its very early stages and we have found its performance to be lagging significantly below the other TSFMs we have considered. We therefore did not include it in the comparison below.
@@
(Rasul et al., 2024) LagLlama is a probabilistic foundation model that extends the Llama 2 architecture (Touvron et al., 2023) to time series data. LagLlama follows a z-score normalization at a window level, and tokenizes the input time series by extracting lagged features from the past values of the time series at different time lags. This allows the model to learn how past observations influence future ones. LagLlama is pretrained on 27 publicly available time series datasets from various domains. The model has about $\sim 2.5$M parameters.
@@
Table 1 lists a few additional details on the various TSLMs described in this section. In particular, it provides information on when the model was released, i.e. when the pretraining sample ends, the list of all the data domains used in the pretraining phase of the model, and the size of training sample. It also flags whether each TSLM is univariate or multivariate.
@@
Table 1: Training datasets for the various TSLMs considered in this paper. For each model, we provide a summary of when the model was pretrained, the list of all the datasets used in the pretraining phase of the model, the size of training sample, and whether it is multivariate. Additional details and references on the various datasets can be found in the individual TSLMs’ technical papers. indicates that the training data include the Monash forecast repository (available at https://forecastingdata.org), and therefore includes a large part of the FRED-MD dataset.